OpenStudy (anonymous):

Solve the differential equation dy/dx=14x^2y^2 with the condition y(2)=4 the solution to the equation is y=

OpenStudy (amistre64):

when given a seperable equation .... seperate it

OpenStudy (anonymous):

how do we separate it?

OpenStudy (amistre64):

\[\frac {dy}{dx}=f(x)~g(y)\] \[\frac {dy}{g(y)}=f(x)~dx\]

OpenStudy (amistre64):

its basic algebra to seperate it

OpenStudy (anonymous):

so itll by dy/y^2=14x^2

OpenStudy (anonymous):

**14x^2dx

OpenStudy (amistre64):

of course: or y^(-2) dy = 14x^2 dx

OpenStudy (anonymous):

separate the variables and integrate. substitute x=2,y=4 and find the value of constant and then substitute the value of constant.

OpenStudy (anonymous):

im lost..

OpenStudy (amistre64):

i cant see your work to determine why you are lost

OpenStudy (anonymous):

or we suppose to find the derrivitve?

OpenStudy (amistre64):

no, the integration

OpenStudy (amistre64):

its already in derivative form

OpenStudy (anonymous):

so itll be y^-1/-1 dy=14x^3/3 dx?

OpenStudy (amistre64):

the dy dx parts are actually part of the operator and disappear dont forget a +C at the end

OpenStudy (anonymous):

so y^-1/-1=14x^3/3+c

OpenStudy (amistre64):

good, now x=2 and y=4, solve for c

OpenStudy (anonymous):

c=-451/12?

OpenStudy (amistre64):

\[\frac 1y=\frac{14}3x^3+C\] \[\frac 14=\frac{14}32^3+C\] \[\frac 14-\frac{14}32^3=C\]

OpenStudy (amistre64):

then rewrite for y in a more normal format of y= .....

OpenStudy (amistre64):

-445/12 i believe

OpenStudy (amistre64):

opps, i forgot the negative onthe y part

OpenStudy (amistre64):

your right

OpenStudy (anonymous):

how do we rewrite it?

OpenStudy (amistre64):

1/y flips to y just place a 1 over the right side to flip it

OpenStudy (amistre64):

\[-\frac1y=\frac{14}3x^3-\frac{451}{12}\] \[\frac1y=\frac{451}{12}-\frac{14}3x^3\] \[y=\frac{1}{\frac{451}{12}-\frac{14}3x^3}\] and simplify to your hearts content

OpenStudy (anonymous):

so what do we do next

OpenStudy (amistre64):

thats it, youve found "y" that fits the information

OpenStudy (anonymous):

thanks! can you help me with another problem similar to this

OpenStudy (anonymous):

dy/dx=4x^4y^4 Y(2)=2

OpenStudy (amistre64):

its not similar, its the exact same thing ...

OpenStudy (amistre64):

seperate, integrate, and solve for C